OpenStudy (anonymous):

solve the system of three equations 2x-3y+5z=13 -3x+3y-2z=-8 5x-2y+4z=2 solve with full steps and work thank you and for every variable using eather substitution method or elimination

OpenStudy (ash2326):

The system of three equations 2x-3y+5z=13 equation 1 -3x+3y-2z=-8 equation 2 5x-2y+4z=2 equation 3 we'll use elimintaion method as well as substitution add the first two equation y will get get cancelled 2x-3y+5z+-3x+3y-2z=13-8 we get -x+3z=5 let this be equation 4 from second equation we have 3y=3x+2z-8 y=(3x+2z-8)/3 substituting this in equation 3 5x-2(3x+2z-8)/3+4z=2 15x-6x-4z+16+12z=6 9x+8z=-10 let this be equation 5 from equation 4 we have x=3z-5 put this in equation 5 27z-45+8z=-10 35z=35 z=1 now x=3z-5 x=3*1-5 x=-2 we had found earlier y=(3x+2z-8)/3 x=-2 and z=1 y=(3*-2+2*1-8)/3 y=(-6+2-8)/3 y=-12/3 y=-4 so x=-2 y=-4 and z=1